solve for the pH of a solution containing 7.5g of ammonium hydroxide in 1.5L soln. Kb of ammonium hydroxide is 4.8x10^-11
1
Expert's answer
2020-10-28T08:41:10-0400

M(NH4OH) = 35 g/mol

n(NH4OH) = m/M = 7.5 g / 35 g/mol = 0.214 mol

C(NH4OH) = n/V = 0.214 mol / 1.5 L = 0.143 mol/L

NH4OH = NH4+ + OH-

0.143 M 0 0

-x +x +x

0.143-x x x

Kb = 4.8x10-11 = (x)(x) / 0.143 - x = x2 / 0.143

x = 2.6x10-6 M

pH = -lg [OH-] = -lg (2.6x10-6) = 5.585